OpenStudy (anonymous):

If f(x) = 2x + 3 and g(x) = 4x – 5, what is f(x) ◦ g(x)?

OpenStudy (amistre64):

is that an open circle?

OpenStudy (amistre64):

if so then it represents a composite; so just stuff g into the x of f

hero (hero):

@amistre64, your explanations are amusing.

hero (hero):

By definition, f(x) ◦ g(x) = f(g(x)) Thus f(g(x)) = 2(4x - 5) + 3 = 8x - 10 + 3 = 8x - 7

OpenStudy (anonymous):

@Hero, thnx, bro

hero (hero):

That is how you "stuff g into the x of f".

OpenStudy (anonymous):

@Hero what does "◦" mean?

hero (hero):

Hang on, I want to make sure I'm saying it right

OpenStudy (amistre64):

lol, thnx

hero (hero):

Apparently, it means "composed with"

hero (hero):

So you would read f(x) ◦ g(x) as "f(x) composed with g(x)"